Circular No. 7357

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/2000 C1 (HERGENROTHER) Carl Hergenrother, Lunar and Planetary Laboratory, reported a new

object (discovery observation below) that showed an 11′′ tail in p.a. 300^{o} on one of four CCD images taken with the 0.41-m Schmidt telescope at Catalina. Following posting on The NEO Confirmation Page, numerous CCD observers have reported cometary appearance, as follows: Feb. 5.3 UT, coma diameter ∼ 12′′, brighter 60′′ tail in p.a. 290^{o}, extending more faintly to 180′′ (J. E. McGaha, Tucson, AZ, 0.62-m reflector); Feb. 5.5, tail ∼ 12′′ long toward the northwest (G. Billings, Calgary, AB, 0.36-m reflector); Feb. 5.7, slightly diffuse with very faint tail ∼ 10′′ long to the northwest (G. J. Garradd, Loomberah, N.S.W., 0.45-m reflector); Feb. 6.1, coma diameter 0′.1, tail 0′.3 long in p.a. 290^{o} (P. Pravec and P. Kušnirák, Ondřejov, 0.65-m reflector); Feb. 6.4, faint tail < 10′′ long in p.a. ∼ 290^{o} (D. T. Durig, Sewanne, TN). Prediscovery observations by LINEAR on Jan. 4 and 8 have also been identified. Complete astrometry and the following orbital elements appear on MPEC 2000-C30:

2000 UT  α_{2000}  δ_{2000}  m_{1}
Feb. 4.45802  12^{h}10^{m}38.^{s}72  +4^{o}53′ 6′′.1  17.1
T = 2000 Mar. 18.7279 TT  ω = 50.^{o}6637
e = 0.410396  Ω = 127.1207 }2000.0
q = 2.099992 AU  i = 6.1192
a = 3.561702 AU n^{o} = 0.1466283 P = 6.722 years

SUPERNOVAE 1999ax AND 1999ay A. Gal-Yam and D. Maoz, Tel Aviv University, write: “A spectrum

(range 440–570 nm) of SN 1999ax (IAUC 7130), obtained by R. A. Stathakis on 1999 Apr. 9 with the 4-m Anglo-Australian Telescope (+ RGO spectrograph), is similar to that of a type-Ia supernova, ∼ 3 weeks after maximum. The redshift of SN 1999ax, estimated from its broad features, is z ∼ 0.05, making it a foreground object to the cluster Abell 1852. Another spectrum (range 340–1000 nm) of SN 1999ax, obtained by G. Aldering on 1999 Apr. 7 with the Kitt Peak 4-m Mayall telescope, confirms the above type and redshift, including a clear detection of the 615-nm Si II feature. Aldering also kindly provided us with a spectrum of SN 1999ay (IAUC 7130), obtained at the same time, showing SN 1999ay to be of type II, at redshift ∼ 0.04, which makes this a foreground object to Abell 1966.”

2000 February 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green